Please help.
Use the distributive property to remove the parentheses in the following expression. Then simplify your result if possible.
1/4(20+16)
Thanks
JMS

Answer by ptaylor(2048)

Distributive property sez:
a(b+c)=ab+ac So:
1/4(20+16)=1/4(20)+1/4(16)=5+4=9
CK
1/4(20+16)=1/4(36)=9
Hope this helps---ptaylor